OpenStudy (anonymous):

the perimeter of the rectangular playing field is 272 yards. the length of the feild is 5 yards less the double the width. what are the dimensions of the playing field?

OpenStudy (anonymous):

w=47 l=89

OpenStudy (anonymous):

 so if the width is \[w\] than the length is \[2w-5\] perimeter is the sides added together so 2(2w-5)+2(w)=272 4w-10+2w=272 6w-10=272 6w=282 w=47 so the width is 47 and the length is 89